OpenStudy (anonymous):

Checking my answers Tell whether the sequence is arithmetic. If it is, what is the common difference? –12, –7, –2, 3, . . . I got no it is not arithmetic

OpenStudy (hba):

Do you know the formula to take the common difference ?

OpenStudy (anonymous):

I just thought you subtracted one from another and if it had a common difference like 4 between each then it was arithmetic

OpenStudy (hba):

If, \[\huge\ T_3-T_2=T_2-T_1\] Then it is arithmetic Here T_1=first term=–12 T_2=second term=–7 T_3=third term= –2 Start working.

OpenStudy (anonymous):

so no its not

OpenStudy (anonymous):

because 5 does not equal -5

OpenStudy (hba):

@ILoVeMaTh3 Please show me your working. You are misleading.

OpenStudy (anonymous):

-2-7=5 -12-(-7)=-5

OpenStudy (hba):

(-)(-)=+

OpenStudy (anonymous):

t2 - t2 = t2-t1 and they do not equal

OpenStudy (anonymous):

*t3

OpenStudy (anonymous):

Am I not allowed to check my answers that I have already gotten and work them out if they are wrong?

OpenStudy (hba):

@ILoVeMaTh3 you are doing it wrong See your terms You are placing them wrong.

OpenStudy (anonymous):

t3 = -2 t2 = -7

OpenStudy (anonymous):

-2-(-7)=5 -7-(-12)=5 Ah so it is arithmetic and 5 is the answer

OpenStudy (anonymous):

I was making -12 positive

OpenStudy (anonymous):

But am I not allowed to check if my answers are correct and work them out if they are wrong with help from people